OpenStudy (anonymous):

Express the complex number in trigonometric form. -4i

OpenStudy (jdoe0001):

keep in mind that \(\bf a+bi \implies r(cos(\theta)+isin(\theta)\\ \bf -4i \implies 0-4i\\ \begin{array}{cccl} x&yi\\ a&bi\\ \hline\\ \bf0&\bf -4i\\ \hline\\ r = \sqrt{a^2+b^2} \end{array}\)

OpenStudy (jdoe0001):

so to convert it, find "r" by using the "a" and "b" components then find the angle by using the "tangent" function with the "a" and "b" components

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."